OpenStudy (anonymous):

Sorry if this is hard to understand. Fractions are hard to write. ((4)/(3a-5))+((3a^2+38a+36)/(25a-9a^3))=((3)/(3a+5))

OpenStudy (danjs):

ok

OpenStudy (danjs):

\[\frac{ 4 }{ 3a-5 } + \frac{ 3a^2+38a+36 }{ 25a-9a^2 } = \frac{ 3 }{ 3a+5 }\]

OpenStudy (danjs):

did i get that right?

OpenStudy (aaliyahfairgood):

Yeah

OpenStudy (danjs):

ok

OpenStudy (anonymous):

yeah that looks right

OpenStudy (danjs):

i would multiply everything by a common denominator , so you can get rid of the fractions

OpenStudy (danjs):

oh thats a 9a^3 in the bottom right?

OpenStudy (anonymous):

Yeah. its cubed. thats also part of the problem why i cant figure it out

OpenStudy (danjs):

\[25a-9a^3 = -a(9a^2 -25) = -a(3a-5)(3a+5)\]

OpenStudy (aaliyahfairgood):

Wouldn't you factor all of them

OpenStudy (danjs):

sometimes problems from books and homework give you clues, so i would have tried the other denominators (3a-5)(3a+5) as factors of the cubed term first

OpenStudy (danjs):

so multiply everything by -a(3a-5)(3a+5)

OpenStudy (anonymous):

ok hang on one sec

OpenStudy (danjs):

ill starty typing it out

OpenStudy (anonymous):

wait i should each fraction by the thing above or each denominator?

OpenStudy (danjs):

multiply both sides by a ratio of 1 \[\frac{ -a(3a-5)(3a+5) }{ -a(3a-5)(3a+5) }\]

OpenStudy (danjs):

all the fractions will be over the common denominator after that ,,

OpenStudy (anonymous):

oh ok thanks

OpenStudy (danjs):

ill start typing it up

OpenStudy (danjs):

\[\frac{ -4a(3a+5) }{ -a(3a-5)(3a+5) } + \frac{ (3a^2+38a+36) }{ -a(3a-5)(3a+5) } = \frac{ -3a(3a-5) }{ -a(3a-5)(3a+5) }\]

OpenStudy (danjs):

see if you cancel out terms in the numerators and denominators, you are back to the starting problem

OpenStudy (danjs):

so now that everything is over a common denominator, you can equate the numerators, same as if you multiplied both sides by -a(3a-5)(3a+5) to cancel the denominators,

OpenStudy (danjs):

\[-4a(3a+5) + (3a^2+38a+36) = -3a(3a-5)\]

OpenStudy (danjs):

multiply out the terms, and combine like powers

OpenStudy (danjs):

\[-12a^2 - 20a +3a^2+38a+36 = -9a^2+15a\]

OpenStudy (danjs):

u with me ?

OpenStudy (anonymous):

Yeah I follow.

OpenStudy (danjs):

see the a^2 terms cancel , you have -9a^2 on both ends

OpenStudy (anonymous):

you left with 18a+36=15a which leads to 3a=-36 which means a = 12

OpenStudy (danjs):

-12

OpenStudy (anonymous):

i mean negative 12

OpenStudy (danjs):

ill do a quick chek on my calc to see if any mistakes were made.. one sec

OpenStudy (anonymous):

I think that's right because -12 is answer b

OpenStudy (danjs):

yep a=-12 is it

OpenStudy (anonymous):

Thanks!

OpenStudy (danjs):

welcome, anytime

OpenStudy (danjs):

just remember if it is a prob from a book or test or something, and you need to factor to get a common denominator, look at the other denominators as they are probably the factors..

OpenStudy (danjs):

like the (3a-5)(3a+5) in this prob

OpenStudy (anonymous):

ok got it